Weighted average (missing value).Solve the equation for x.

30(2)+40(3)+36(3) +5x/10=34.8
1
Expert's answer
2020-10-22T17:33:10-0400

(30(2)+40(3)+36(3) +5x)/10=34.8

60+120+108 +5x=34.8*10=348

5x=348-60-120-108 =60

x=60/5=12